OpenStudy (anonymous):

If 20 bakers can bake 40 pizzas in 8 hours, how many bakers can bake 10 pizzas in 2 hours? the solution is to get the rate (in man-hour) of baking pizza.

OpenStudy (anonymous):

word problem

hartnn (hartnn):

20 bakers can bake 40 pizzas in 8 hours, so how many pizzas will they(20 bakers) make in 1 hour ?

hartnn (hartnn):

@katrina_gay09 can u answer that ^^

OpenStudy (anonymous):

@katrina_gay09 use this formula N1 * T1 * W2 =N2*T2*W1 where , N1=No of Persons T1=Time taken W1=work N2=No of Persons T2=Time taken W2=work